OpenStudy (anonymous):

solve -4(x+4)=16 a) 0 b) -3 c) -8 d) 8 solve b=-4(c-2d) for d y varies indirectly as twice x. when x=9, y=2.Find y when x=3. a) 9 b) 12 c) 4 d) 6 what is the solution set for |2x-3|=11 a)-4 and 4 b) 7 and -7 c) 7 and -4 d) -7 and 4

OpenStudy (amistre64):

divide out the -4 then subtract 4

OpenStudy (anonymous):

ok so the answer is?

OpenStudy (anonymous):

-8

OpenStudy (anonymous):

b=-4(c-2d) b/-4=c-2d 2d=c+b/4 d=c/2+b/8

OpenStudy (amistre64):

id say the answer is to do your own work ...

OpenStudy (anonymous):

y=2kx 2=18k k=1/9 y=2x/9

OpenStudy (anonymous):

2x-3=+/-11 solve for x

OpenStudy (kirbykirby):

And just a tip if you're going to have multiple choice exams with "solve this",and don't know how to do it, just plug in the numbers from every answer until the equation is satisfied